Explain how the graph of f can be obtain from the graph y=1/x .

f(x)=(-1/x-2)-2

the graph can be obtain from y=1/x by reflecting across the x-axis , and shifting it how?
Answer #1
step 1: draw graph of y=1/x
step 2: shift graph to the right by 2 units to get the graph of y=1/(x-2)
step3: now take reflection of the graph about y=0 to get y=-1/(x-2)
step4:shift the graph upward by 2 units to get graph of y=(-1/(x-2))-2
answered by: waseem
Answer #2

uploaded image

graph of 1/x-2 is

uploaded image

now we will draw the graph of -1/x-2

uploaded imagenowthe graph of 1/(x-2)-2

uploaded image

answered by: Alisa